What is the distance between the points x and y, given that x = 3 + 2i and y = 4 + 3i?
1
Expert's answer
2019-04-15T11:01:06-0400
"|x-y|=\sqrt{(x_1 - y_1)^2+(x_2 - y_2)^2}="

"=\sqrt{(3-4)^2+(2-3)^2}=\sqrt{2}"